21- Seaside issues a bond that has a stated interest rate of 8%, face amount of $50,000, and is due in 6 years. Interest payments are made semi-annually. The market rate for this type of bond is 10%. What is the issue price of the bond? (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.)

Multiple Choice

  • $45,569.

  • $82,802.

  • $77,842.

  • $50,000.
Answer #1
Bond characterstics Amount
Principal 50,000
interest                2,000
Market interest rate 5%
periods to maturity 12
issue price 45,569 Answer
Calculation of bond issue price
Where
i= 5.00%
t= 12
principal * PV of $1 at 5% for 12 yrs =
50,000 * 0.55684        = 27842
interest * PV of ordinary annuity at 5%=
2000 * 8.86325 = 17727
bond issue price 45569